OpenStudy (anonymous):

Using the Half Angle Identity find sin13pi/12 (find exact value)

OpenStudy (anonymous):

@iGreen plz help!!!

OpenStudy (adamaero):

So are you suppose to split this up, to make half angle identities? If so, show me what ya got

OpenStudy (anonymous):

alright @adamaero here is what I got sin(13pi/12) --> (pi+pi/12) --> -sin(pi/12) --> sqrt(1-cos(pi/6)/2) --> sqrt(1-cos[sqrt(3/2)]/2) --> sqrt(2-sqrt(3/4)) --> sqrt(2-sqrt(3/2))

OpenStudy (adamaero):

wouw, just press enter gtg @xapproachesinfinity @Luigi0210 @zepdrix

OpenStudy (anonymous):

sorry that I made so messy so here is my finaql answer... \[\sqrt{2-\sqrt{3}}/2 \]

OpenStudy (anonymous):

what do ya think xapproach?

OpenStudy (anonymous):

You must have a lot to say :o

OpenStudy (xapproachesinfinity):

what you did was not half angle identity \[\sin^2(\frac{13\pi}{12})=\frac{1-\cos (\frac{13\pi}{6})}{2}\]

OpenStudy (xapproachesinfinity):

and 13pi/6 can be written as 2pi+pi/6 and since cos has 2pi period then cos (2pi+pi/6)=cos (pi/6)

OpenStudy (xapproachesinfinity):

oh i was away actually doing something else lol

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

we both got to cos(pi/6) though

OpenStudy (xapproachesinfinity):

this is the half angle identity : \[\sin^2\theta=\frac{1-\cos 2\theta}{2}\]

OpenStudy (xapproachesinfinity):

well you definitely get the same answer with other methods they all work fine

OpenStudy (xapproachesinfinity):

just the question asked you to do it with half angle identity for anything that gets you to the answer is legitimate

OpenStudy (anonymous):

alrighty thanks!

OpenStudy (xapproachesinfinity):

YW